# AdaptInfer: Adaptive Token Pruning for Vision–Language Model Inference with Dynamical Text Guidance

### v1

@@ Abstract
Vision–language models (VLMs) have achieved impressive performance on multimodal reasoning tasks such as visual question answering (VQA), but their inference cost remains a significant challenge due to the large number of vision tokens processed during the prefill stage. Existing pruning methods often rely on directly using the attention patterns or static text prompt guidance, failing to exploit the dynamic internal signals generated during inference. To address these issues, we propose AdaptInfer, a plug-and-play framework for adaptive vision token pruning in VLMs. First, we introduce a fine-grained, dynamic text-guided pruning mechanism that reuses layer-wise text-to-text attention maps to construct soft priors over text-token importance, allowing more informed scoring of vision tokens at each stage. Second, we perform an offline analysis of cross-modal attention shifts and identify consistent inflection locations in inference, which inspire us to propose a more principled and efficient pruning schedule. Our method is lightweight and plug-and-play, also generalizable across multi-modal tasks. Experimental results have verified the effectiveness of the proposed method. For example, it reduces CUDA latency by 61.3% while maintaining an average accuracy of 92.9% on vanilla LLaVA-1.5-7B. Under the same token budget, AdaptInfer surpasses SOTA in accuracy.
Introduction
@@
In recent years, building on the success of LLMs (Bommasani et al. 2021; Touvron et al. 2023; Brown et al. 2020), vision–language models (VLMs) have emerged to tackle multimodal reasoning by combining visual encoders ((Liu et al. 2022; Dosovitskiy et al. 2021)), with LLMs’ text decoders (Du et al. 2022). This integration enables impressive performance on tasks such as captioning (Lin et al. 2014), image retrieval (Faghri et al. 2018), and visual question answering (VQA) (Antol et al. 2015), but it also introduces a new computational challenge: the sheer number of vision tokens.
@@
During the inference process of VLMs, the number of vision tokens is often much larger than that of textual tokens, sometimes by an order of magnitude or more. For instance, an image of size $672\times 672$ processed by a visual encoder with a patch size of $14\times 14$ typically results in $2304$ vision tokens (Radford et al. 2021; Zhang et al. 2025), whereas the corresponding text prompt may contain fewer than 100 tokens (Hudson and Manning 2019; Fu et al. 2023). Also, much previous research suggests that the vision tokens are more redundant and semantically repetitive (Zhang et al. 2024; Tong et al. 2025).
@@
As a result, explorations in VLM acceleration primarily focus on the efficient pruning, compression or sparsification of vision tokens. This paradigm aims to reduce computational overhead by retaining only the most valuable vision tokens. Among them, some works introduce sparsity strategies within the visual encoders to generate less but useful enough vision tokens (Li, Wang, and Jia 2024; Chen et al. 2025b), while others further prune tokens based on either self-attention or cross-attention patterns during the prefill stage (Lin et al. 2025; Xing et al. 2024; Chen et al. 2025a; Li et al. 2025). Nevertheless, not all the attention logits should be involved in the vision token ranking for the dispersion of the full attention patterns (Zhang et al. 2024). Granting voting rights only to the most salient tokens sharpens guidance, enabling more aggressive yet accurate vision‑token pruning.
@@
To address this, SparseVLM (Zhang et al. 2025) introduces the concept of text prompt–guided pruning, selecting the most salient text tokens offline before the prefill pass. While this approach acknowledges the importance of textual cues, it does not fully address the underlying challenge: the dynamic nature of text token importance during inference. In practice, the informativeness of text tokens evolves across layers as the model progressively refines its internal representations (Tenney, Das, and Pavlick 2019; Clark et al. 2019). Our observations in Figure 2 also indicate that the most prominent text tokens vary significantly across layers, making any static selection inherently suboptimal.
Figure 1: Illustration of Dynamic Text‑Guided Attention in VLMs. A schematic illustration showing the set of key text tokens and the corresponding visual regions shift across transformer layers.
@@
Moreover, current methods determine pruning hyperparameters (e.g. the pruning locations) primarily through either empirical rule-of-thumb or extensive hyperparameter optimization experiments (Xing et al. 2024; Chen et al. 2025a; Zhang et al. 2024). However, we argue that relying on manual tuning or grid search not only imposes substantial offline computational overhead, but also leads to task- or dataset-specific heuristics. In this work, we take the first step in providing a principled pruning schedule. We provide our insights based on systematically analyzing the distributional characteristics of attention shifts to vision tokens during VLM inference. Specifically, we identify consistent attention inflection points at layer 1, 10, and 20 in LLava-1.5-7B (Liu et al. 2023b), suggesting that aggressive pruning immediately after these layers is a more effective and computationally efficient strategy on LLava.
@@
Vision-Language Models
Early multi-modal systems paired convolutional vision backbones with recurrent language decoders (Karpathy and Fei-Fei 2017; Vinyals et al. 2015). Modern VLMs instead follow the Transformer paradigm (Vaswani et al. 2017) by representing an image as a sequence of visual tokens that interact with textual tokens in a shared self-attention space (Liu et al. 2023b; Chen et al. 2024a, b).
@@
BLIP-2 (Li et al. 2023) and MiniGPT-4 (Zhu et al. 2023) introduce lightweight linear adapters that project features from a frozen CLIP encoder into the hidden space of a large language model, enabling efficient training. These explorations bridge the gap between frozen encoders and LLMs. The LLaVA family (Liu et al. 2023a, b, 2024a) refines this recipe with stronger instruction tuning, while a few other efforts such as Flamingo (Alayrac et al. 2022), CogVLM (Wang et al. 2024) and GPT-4V (OpenAI 2023) scale the approach to billions of parameters.
@@
Recent explorations also contribute to resolution and multi-modal extensions. Models leveraging hierarchical perception (e.g., LLaVA‑NeXT (Liu et al. 2024a)) and adaptive patching (e.g., Qwen‑VL (Bai et al. 2023)) permit high‑resolution inputs, while large-scale vision encoders are adopted to generate richer hidden states of the vision tokens (Chen et al. 2024a, b). However, these gains come at the cost of dramatically more vision tokens, which is a key bottleneck addressed by our work.
Inference Acceleration of VLMs
@@
Previous approaches mainly focus on vision token sparsification. This is because the number of vision tokens is often an order of magnitude (or more) larger than that of text tokens. In addition, visual embeddings are naturally much more sparse and repetitive than human-made texts (Marr 2010). In this field, there are two research directions including efficient vision encoders and vision token pruning in LLM networks.
@@
For example, methods like LLaVA-PruMerge (Shang et al. 2025) and FlowCut (Tong et al. 2025) follow the first direction, which cuts the encoder outputs or uses a lightweight projector to reduce the number of vision tokens. Recoverable compression (Chen et al. 2025b) is also introduced to repair the information loss from token pruning within vision encoders. Solutions follow the second direction not only to drop the vision tokens (Chen et al. 2025a; Lin et al. 2025; Xing et al. 2024; Li et al. 2025; Luan et al. 2025; Ye et al. 2025) during prefill, but also merge them to compress the numbers of the vision tokens (Bolya et al. 2023) and recover them in certain inference stage (Wu, Chen, and Wang 2025). SparseVLM (Zhang et al. 2025) tries to go deeper in exploration of static text prompt guidance but ignoring the evolving inherent of token information.
Method
Figure 2: MIoU of key text tokens in different layers. Most values are lower than 0.5, which indicates the text token importance varies significantly.
(a) Attention shift points distribution on MME.
(b) Attention shift points distribution on TextVQA.
Figure 3: Layer-wise distribution of attention shifts. The shift points appears consistently on two datasets. Results are aggregated over 1,000 samples for each dataset.
@@
Observations
Text Token Importance
Given an input question prompt, we first tokenize it as $Prompt=[t_{1},\ldots,t_{n}]$. We then define the importance of a text token $t_{i}$ at layer $\ell$ as the total attention weight it receives from all text tokens in the layer-$\ell$ text-to-text (t2t) attention map:
$\mathrm{Imp}_{\ell}(t_{i})=\sum_{j=1}^{n}\mathbf{A}^{(\ell)}_{\text{t2t}}[j,i],$ (1)
where $\mathbf{A}^{(\ell)}_{\text{t2t}}[j,i]$ denotes the attention directed from token $t_{j}$ to $t_{i}$. Intuitively, tokens with high $\mathrm{Imp}_{\ell}$ are the current key text tokens of the language stream and are therefore best suited to guide cross-modal pruning decisions at that layer.
Dynamics of Text Token Importance
To demonstrate that the importance of text tokens evolves significantly across layers, we conduct a simple empirical study on LLava-1.5-7B (Liu et al. 2023b). LLava-1.5-7B contains a 32-layer LLaMa (Touvron et al. 2023) as its language model. We extract t2t attention maps from 1,000 samples in the TextVQA dataset. On average, each sample contains approximately 100 text tokens. At each chosen layer, we select the top 20% text tokens as the key text tokens of that layer. We then compute the mean Intersection over Union (mIoU) between the indexes of the selected top tokens across layers, as reported in Figure 2. The value in the i-th row and j-th column represents the mIoU of the key text token indexes of the i-th and j-th layers. Note that this figure is symmetrical, and the miou values along the diagonal are always 1 because they are comparisons within the same layer.
Cross-Attention Shifts of VLM
We argue that a principled approach to setting pruning hyperparameters is not only necessary but preferable to complex, trial-and-error–based tuning. To support this claim, we performed an analysis to investigate the locations of cross-attention shifts during inference.
Specifically, we calculate cumulative text-to-vision (t2v) attention scores for visual tokens at each layer using LLava-1.5-7B (Liu et al. 2023b). For each sample, we first select the top 10% of vision tokens, approximately 58 tokens per image, which receive the highest total attentions in the prefill stage. Similar to the text tokens, these vision tokens are assumed to be the most critical ones for the corresponding image-based VQA task.
To understand when these important tokens become semantically salient, we analyze their cumulative attention trajectories across transformer layers. We apply change-point detection (Truong, Oudre, and Vayatis 2020) on each curve to identify the layer where the model’s attention pattern changes significantly. The technical details of change-point detection will be presented in the appendix. Intuitively, an attention shift point may indicate that either (1) the token begins to receive significantly more attention, becoming critical, or (2) the token’s informative content has already been fully extracted, becoming redundant. In both cases, these shifts mark the layer-wise transitions in how the model utilizes visual information.
Figure 3 shows the distribution of the attention shift locations aggregated from 1,000 samples each in the MME and TextVQA datasets. Despite data set differences, we observe a highly consistent trend: attention shifts cluster densely in layer 1 and round layers 10-20, while layers 2-8 and 20 + show a frequency of near-zero attention shifts.
These findings provide a data-driven basis for pruning schedule design, where pruning locations are informed by the model’s own attention behavior, rather than by empirical intuition alone.
Adaptive Token Pruning
@@
Dynamic Text Guidance Pruning
To guide visual token pruning in a more informed and adaptive way, we propose a dynamic text-guidance mechanism that leverages the model’s internal attention signals. Instead of statically selecting a fixed subset of text tokens before entering the language model (Zhang et al. 2025), we dynamically infer their relative importance during inference at each predefined pruning layer. We reuse the attention maps already computed by the model. The architecture of Our AdaptInfer is shown in Fig 4, and our dynamic text guidance mechanism can be divided into three steps.
Firstly, at each pruning layer, we extract the text-to-text attention matrix $\mathbf{A}_{\text{t2t}}\in\mathbb{R}^{T\times T}$, where $T$ is the number of text tokens. To estimate the importance of each text token, we aggregate attention scores across the query dimension:
$\mathbf{w}=\sum_{i=1}^{T}\mathbf{A}_{\text{t2t}}[i,:]\in\mathbb{R}^{T},$ (2)
where $\mathbf{w}$ serves as a soft prior distribution over the text tokens, indicating how much attention each token receives from the rest of the sequence.
Secondly, we then use this prior to reweight the t2v attention matrix $\mathbf{A}_{\text{t2v}}\in\mathbb{R}^{T\times V}$ at the same layer, where $V$ is the number of visual tokens remained. $\mathbf{A}_{\text{t2v}}$ denotes the cross-attention matrix where text tokens are used as queries, and vision tokens serve as keys and values. The importance score of each visual token is computed as:
$\mathbf{s}=\mathbf{w}^{\top}\cdot\mathbf{A}_{\text{t2v}}\in\mathbb{R}^{V}.$ (3)
@@
Analysis of Computational Complexity

### v2

@@ Abstract
Vision–language models (VLMs) have achieved impressive performance on multimodal reasoning tasks such as visual question answering, image captioning and so on, but their inference cost remains a significant challenge due to the large number of vision tokens processed during the prefill stage. Existing pruning methods often rely on directly using the attention patterns or static text prompt guidance, failing to exploit the dynamic internal signals generated during inference. To address these issues, we propose AdaptInfer, a plug-and-play framework for adaptive vision token pruning in VLMs. First, we introduce a fine-grained, dynamic text-guided pruning mechanism that reuses layer-wise text-to-text attention maps to construct soft priors over text-token importance, allowing more informed scoring of vision tokens at each stage. Second, we perform an offline analysis of cross-modal attention shifts and identify consistent inflection locations in inference, which inspire us to propose a more principled and efficient pruning schedule. Our method is lightweight and plug-and-play, also generalizable across multi-modal tasks. Experimental results have verified the effectiveness of the proposed method. For example, it reduces CUDA latency by 61.3% while maintaining an average accuracy of 93.1% on vanilla LLaVA-1.5-7B. Under the same token budget, AdaptInfer surpasses SOTA in accuracy.
1 Introduction
@@
In recent years, building on the success of LLMs (Bommasani et al., 2021; Touvron et al., 2023; Brown et al., 2020), vision–language models (VLMs) have emerged to tackle multimodal reasoning by combining visual encoders  (Liu et al., 2022; Dosovitskiy et al., 2021), with LLMs’ text decoders (Du et al., 2022). This integration enables impressive performance on tasks such as captioning (Lin et al., 2014), image retrieval (Faghri et al., 2018), and visual question answering (VQA) (Antol et al., 2015), but it also introduces a new computational challenge: the sheer number of vision tokens.
@@
During the inference process of VLMs, the number of vision tokens is often much larger than that of textual tokens, sometimes by an order of magnitude or more. For instance, an image of size $672\times 672$ processed by a visual encoder with a patch size of $14\times 14$ typically results in $2304$ vision tokens (Radford et al., 2021; Zhang et al., 2025), whereas the corresponding text prompt may contain fewer than 100 tokens (Hudson & Manning, 2019; Fu et al., 2023). Also, much previous research suggests that the vision tokens are more redundant and semantically repetitive (Zhang et al., 2024; Tong et al., 2025).
@@
As a result, explorations in VLM acceleration primarily focus on the efficient pruning, compression or sparsification of vision tokens. This paradigm aims to reduce computational overhead by retaining only the most valuable vision tokens. Among them, some works introduce sparsity strategies within the visual encoders to generate less but useful enough vision tokens (Li et al., 2024; Chen et al., 2025b), while others further prune tokens based on either self-attention or cross-attention patterns during the prefill stage (Lin et al., 2025; Xing et al., 2024; Chen et al., 2025a; Li et al., 2025). Nevertheless, not all the attention logits should be involved in the vision token ranking for the dispersion of the full attention patterns (Zhang et al., 2024). Granting voting rights only to the most salient tokens sharpens guidance, enabling more aggressive yet accurate vision‑token pruning.
@@
To address this, SparseVLM (Zhang et al., 2025) introduces the concept of text prompt–guided pruning, selecting the most salient text tokens offline before the prefill pass. While this approach acknowledges the importance of textual cues, it does not fully address the underlying challenge: the dynamic nature of text token importance during inference. In practice, the informativeness of text tokens evolves across layers as the model progressively refines its internal representations (Tenney et al., 2019; Clark et al., 2019). Our observations in Figure 1(a) also indicate that the most prominent text tokens vary significantly across layers, making any static selection inherently suboptimal.
@@
Moreover, current methods determine pruning hyperparameters (e.g. the pruning locations) primarily through either empirical rule-of-thumb or extensive hyperparameter optimization experiments (Xing et al., 2024; Chen et al., 2025a; Zhang et al., 2024). However, we argue that relying on manual tuning or grid search not only imposes substantial offline computational overhead, but also leads to task- or dataset-specific heuristics. In this work, we take the first step in providing a principled pruning schedule. We provide our insights based on systematically analyzing the distributional characteristics of attention shifts to vision tokens during VLM inference. Specifically, we identify consistent attention inflection points at layer 1, 10, and 20 in LLava-1.5-7B (Liu et al., 2023b) and at layer 0, 9, and 19 on Qwen2-VL-2B (Wang et al., 2024a), suggesting that aggressive pruning immediately after these layers is a more effective and computationally efficient strategy on LLava and Qwen.
@@ 2 Related Work
2.1 Vision-Language Models
@@
Early multi-modal systems paired convolutional vision backbones with recurrent language decoders (Karpathy & Fei-Fei, 2017; Vinyals et al., 2015). Modern VLMs instead follow the Transformer paradigm (Vaswani et al., 2017) by representing an image as a sequence of visual tokens that interact with textual tokens in a shared self-attention space (Liu et al., 2023b; Chen et al., 2024a; b). BLIP-2 (Li et al., 2023a) and MiniGPT-4 (Zhu et al., 2023) introduce lightweight linear adapters that project features from a frozen CLIP encoder into the hidden space of a large language model, enabling efficient training. These explorations bridge the gap between frozen encoders and LLMs. The LLaVA family (Liu et al., 2023a; b; 2024a) refines this recipe with stronger instruction tuning, while a few other efforts such as Flamingo (Alayrac et al., 2022), CogVLM (Wang et al., 2024b) and GPT-4V (OpenAI, 2023) scale the approach to billions of parameters.
@@
Recent explorations also contribute to resolution and multi-modal extensions. Models leveraging hierarchical perception (e.g., LLaVA‑NeXT (Liu et al., 2024a)) and adaptive patching (e.g., Qwen‑VL (Bai et al., 2023)) permit high‑resolution inputs, while large-scale vision encoders are adopted to generate richer hidden states of the vision tokens (Chen et al., 2024a; b). However, these gains come at the cost of dramatically more vision tokens, which is a key bottleneck addressed by our work.
2.2 Inference Acceleration of VLMs
@@
Previous approaches mainly focus on vision token sparsification. This is because the number of vision tokens is often an order of magnitude (or more) larger than that of text tokens. In addition, visual embeddings are naturally much more sparse and repetitive than human-made texts (Marr, 2010). In this field, there are two research directions including efficient vision encoders and vision token pruning in LLM networks.
@@
For example, methods like LLaVA-PruMerge (Shang et al., 2025) and FlowCut (Tong et al., 2025) follow the first direction, which cuts the encoder outputs or uses a lightweight projector to reduce the number of vision tokens. Recoverable compression (Chen et al., 2025b) is also introduced to repair the information loss from token pruning within vision encoders. Solutions follow the second direction not only to drop the vision tokens (Chen et al., 2025a; Lin et al., 2025; Xing et al., 2024; Li et al., 2025; Luan et al., 2025; Ye et al., 2025) during prefill, but also merge them to compress the numbers of the vision tokens (Bolya et al., 2023) and recover them in certain inference stage (Wu et al., 2025). SparseVLM (Zhang et al., 2025) tries to go deeper in exploration of static text prompt guidance but ignoring the evolving inherent of token information. Our approach contributes to the second paradigm.
3 Method
(a) MIoU of key text tokens in different layers.
(b) Attention shift distribution on MME.
(c) Attention shift distribution on TextVQA.
Figure 1: Preliminary observations. (a) The consistent low mIoU shows the key text token varies across layers. (b)(c) Layer-wise distribution of attention shifts on MME and TextVQA, which shows a highly consistent trend.
@@ 3 Method
3.1 Observations
3.1.1 Text Token Importance
Given an input question prompt, we first tokenize it as $Prompt=[t_{1},\ldots,t_{n}]$. We then define the importance of a text token $t_{i}$ at layer $\ell$ as the total attention weight it receives from all text tokens in the layer-$\ell$ text-to-text (t2t) attention map on average of attention heads:
$\mathrm{Imp}_{\ell}(t_{i})=\frac{1}{H}\sum_{h=1}^{H}\sum_{j=1}^{n}\mathbf{A}^{(\ell,h)}_{\text{t2t}}[j,i],$ (1)
where $\mathbf{A}^{(\ell,h)}_{\text{t2t}}[j,i]$ denotes the attention of head $h$ directed from token $t_{j}$ to $t_{i}$, $H$ represents the number of the attention heads. Intuitively, tokens with high $\mathrm{Imp}_{\ell}$ are the current key text tokens of the language stream and are therefore best suited to guide cross-modal pruning decisions at that layer.
3.1.2 Dynamics of Text Token Importance
To demonstrate that the importance of text tokens evolves significantly across layers, we conduct a simple empirical study on LLava-1.5-7B (Liu et al., 2023b). LLava-1.5-7B contains a 32-layer LLaMa (Touvron et al., 2023) as its language model. We extract t2t attention maps from 1,000 samples in the TextVQA dataset. On average, each sample contains approximately 100 text tokens. At each chosen layer, we select the top 20% text tokens as the key text tokens of that layer. We then compute the mean Intersection over Union (mIoU) between the indexes of the selected top tokens across layers, as reported in Figure 1(a). The value in the i-th row and j-th column represents the mIoU of the key text token indexes of the i-th and j-th layers. Note that this figure is symmetrical, and the miou values along the diagonal are always 1 because they are comparisons within the same layer.
3.1.3 Cross-Attention Shifts of VLM
We argue that a principled approach to setting pruning hyperparameters is not only necessary but preferable to complex, trial-and-error–based tuning. To support this claim, we performed an analysis to investigate the locations of cross-attention shifts during inference. Specifically, we calculate cumulative text-to-vision (t2v) attention scores for visual tokens at each layer using LLava-1.5-7B (Liu et al., 2023b). For each sample, we first select the top 10% of vision tokens, approximately 58 tokens per image, which receive the highest total attentions in the prefill stage. Similar to the text tokens, these vision tokens are assumed to be the most critical ones for the corresponding image-based multi-modal task.
To understand when these important tokens become semantically salient, we analyze their cumulative attention trajectories across transformer layers. We apply change-point detection (Truong et al., 2020) on each curve to identify the layer where the model’s attention pattern changes significantly. The technical details of change-point detection are presented in the Appendix B. Intuitively, an attention shift point may indicate that either (1) the token begins to receive significantly more attention, becoming critical, or (2) the token’s informative content has already been fully extracted, becoming redundant. In both cases, these shifts mark the layer-wise transitions in how the model utilizes visual information.
Figure 1(b) and 1(c) shows the distribution of the attention shift locations aggregated from 1,000 samples each in the MME (Fu et al., 2023) and TextVQA (Singh et al., 2019) datasets. Despite data set differences, we observe a highly consistent trend: attention shifts cluster densely in layer 1 and round layers 10-20, while layers 2-9 and 20 + show a frequency of low attention shifts. These findings provide a data-driven basis for pruning schedule design, where pruning locations are informed by the model’s own attention behavior, rather than by empirical intuition alone.
3.2 Adaptive Token Pruning
@@
3.2.1 Dynamic Text Guidance Pruning
To guide visual token pruning in a more informed and adaptive way, we propose a dynamic text-guidance mechanism that leverages the model’s internal attention signals. Instead of statically selecting a fixed subset of text tokens before entering the language model (Zhang et al., 2025), we dynamically infer their relative importance during inference at each predefined pruning layer. We reuse the attention maps already computed by the model. The architecture of Our AdaptInfer is shown in Fig 2, and our dynamic text guidance mechanism can be divided into three steps.
Firstly, in each pruning layer, we extract the text-to-text attention matrix $\mathbf{A}^{h}_{\text{t2t}}\in\mathbb{R}^{T\times T}$, where $T$ is the number of text tokens, $h$ is the index of attention head with a total number of $H$. To estimate the importance of each text token, we aggregate attention scores across the query dimension:
$\mathbf{w^{(h)}}=\sum_{i=1}^{T}\mathbf{A}^{(h)}_{\text{t2t}}[i,:]\in\mathbb{R}^{T},$ (2)
where $\mathbf{w}$ serves as a soft prior distribution over the text tokens, averaged across all attention heads, indicating how much attention each token receives from the rest of the sequence.
Secondly, we then use this prior to reweight the t2v attention matrix $\mathbf{A}^{(h)}_{\text{t2v}}\in\mathbb{R}^{T\times V}$, where $V$ is the number of visual tokens remained. $\mathbf{A}^{(h)}_{\text{t2v}}$ denotes the cross-attention matrix where text tokens are used as queries, and vision tokens serve as keys and values. The importance score of each visual token on average of all attention heads is computed as:
$\mathbf{s}=\frac{1}{H}\sum_{h=1}^{H}\mathbf{w}^{(h)\top}\cdot\mathbf{A}^{(h)}_{\text{t2v}}\in\mathbb{R}^{V}.$ (3)
@@ 3.2.1 Dynamic Text Guidance Pruning
3.2.2 Analysis of Computational Complexity